OpenStudy (mendicant_bias):

(ODE) another elementary Differential Equation, prompt placed below shortly.

OpenStudy (mendicant_bias):

http://i.imgur.com/0T9LWyn.png

OpenStudy (mendicant_bias):

Thinking about how to do this, first off.

OpenStudy (mendicant_bias):

Is it maybe separable? Gonna see if I can make it a separable eqn. @Concentrationalizing

OpenStudy (anonymous):

SHould I actually answer that question or let you figure it out? Haha

ganeshie8 (ganeshie8):

it looks homogeneous

OpenStudy (anonymous):

I dont want you to waste your time, so go with bernoulli :P

OpenStudy (mendicant_bias):

Nevermind, I'm back, site loading issues, keep having to freaking clear my browser history every five minutes while using OpenStudy if I expect it to work.

OpenStudy (perl):

divide both sides by x^2 * y^2

OpenStudy (mendicant_bias):

Thank you! Alright, now I need to remember how to solve a Bernoulli DE...lol.

OpenStudy (mendicant_bias):

Alright, going to put it in that general form.

OpenStudy (mendicant_bias):

\[\frac{dy}{dx}+\frac{y^2}{x^2}=\frac{y}{x}\]

OpenStudy (mendicant_bias):

Now some kind of substitution, m8

OpenStudy (anonymous):

Well, you would need the higher power of y on the other side. So \[\frac{ dy }{ dx }-\frac{ y }{ x }= -\frac{ y^{2} }{ x^{2} }\] Do you remember the substitution to make?

OpenStudy (mendicant_bias):

http://i.imgur.com/aR14ol5.png

OpenStudy (mendicant_bias):

No, but I found it, lol

OpenStudy (anonymous):

Cool, cool.

OpenStudy (mendicant_bias):

Alright, one sec, in the process of becoming slowly less incompetent

OpenStudy (perl):

x^2 * dy/dx + y^2 = xy subtract xy and y^2 from both sides x^2 * dy/dx - xy = -y^2 divide both sides by -x^2 * y^2 (dy/dx)/(-y^2) + 1/ ( x*y) = 1/x^2 let v(x) = 1/y dv/dx = -1/y^2 * dy/dx so we have dv/dx + v(x) / x = 1/x^2 this is a linear ode , we can solve by integrating factor

OpenStudy (mendicant_bias):

Do you think Bernoulli or Integrating Factor would be better, and why?

ganeshie8 (ganeshie8):

this can be worked less painfully without bernouli

ganeshie8 (ganeshie8):

notice that the de is homogeneous and simply substitute `y = vx`

ganeshie8 (ganeshie8):

you will get a separable equation straight

OpenStudy (mendicant_bias):

Alright, cool. I need to learn how to do Bernoulli nonetheless, so I may end up doing it that way. It's homogeneous once we subtract xy, correct? Previously it's not homogeneous unless you're talking about some homogeneous function stuff that IDR.

OpenStudy (perl):

yes homogenous is easier approach

ganeshie8 (ganeshie8):

homogeneous is a cool thing to learn too

OpenStudy (anonymous):

I guess the placement of dy and dx doesnt ruin the idea of homogenous. Guess seeing it the way it was didnt make me thing homogenous :/

OpenStudy (mendicant_bias):

Alright, how about we do both? lol. We'll already know the solution at some given point, so reworking it with a different technique won't be bad.

OpenStudy (mendicant_bias):

Okay, so is this a homogeneous function thing? Need to look up/remember how to deal with those. (I know we literally like just did one of those @Concentrationalizing , wasn't this the "of order blahblahblah" thing?)

OpenStudy (anonymous):

Yeah. To be proper, you would show that \(f(tx,ty) = t^{\alpha}f(x,y)\) for some \(\alpha\). But looking at the combined total of powers in each term, you can see this would be homogenous degree 2. The total amount of powers in each term is 2.

OpenStudy (mendicant_bias):

WHOOOO GUESS WHO JUST HAD TO clEAR HIS HISTORY AGAIN YAY OPENSTUDY

OpenStudy (mendicant_bias):

Alright, site is finally back up for me again.

OpenStudy (mendicant_bias):

Alright, so homogeneous function of order 2. Now to do the substitution @Concentrationalizing

OpenStudy (mendicant_bias):

(PS I am still legitimately intent as solving it as a Bernoulli DE, because I need to familiarize myself with it)

OpenStudy (mendicant_bias):

\[y=ux.\]

OpenStudy (mendicant_bias):

(I'm currently looking over Ganeshie and Perl's stuff to catch up to what they did, sorry, one minute)

OpenStudy (mendicant_bias):

Alright, so just getting this clear, when you're evaluating the order of a DE to see whether it is homogeneous or not, derivatives don't contribute at all, right? e.g. y' won't contribute to the degree/order of the homogeneous function.

OpenStudy (mendicant_bias):

But yeah, looks like a homogeneous function of degree 2.

OpenStudy (anonymous):

No, it wont. Just check the total degee of each term. All of them are 2 this time so yeah, degree 2.

OpenStudy (mendicant_bias):

And I'm going to quickly read over the other problem we just did to make sure I remember how to do this, if it isn't clear yet my general memory span is horrific

OpenStudy (mendicant_bias):

In case anybody in the future happens to find this thead through google site search and wants to see a worked Homogeneous function problem I'm mentioning, here you go: http://openstudy.com/study#/updates/54850fbbe4b01e7eabcc0904

OpenStudy (mendicant_bias):

Alright, going to try to get it into the form \[M(x,y)dx+N(x,y)dy=0\]now. I think Ganeshie or Perl already did that, but they didn't do it in LaTeX and it was frankly hard to read.

OpenStudy (mendicant_bias):

\[x^2 \frac{dy}{dx}+y^2=xy\]

OpenStudy (anonymous):

Perl basically went along the lines of bernoulli set-up. So this is completely different.

OpenStudy (mendicant_bias):

Well, one of them did, I think Ganeshie did, oh well, w/e

OpenStudy (mendicant_bias):

Alright, I want to solve it as Bernoulli first anyways, can we do that?

OpenStudy (mendicant_bias):

So, putting it in the appropriate form of a Bernoulli DE:

OpenStudy (anonymous):

He's the one who had the idea, but none of us actually did it, lol. But sure, put it into the form ya need.

OpenStudy (mendicant_bias):

\[\frac{dy}{dx}+\frac{y^2}{x^2}=\frac{y}{x}\]

OpenStudy (mendicant_bias):

Now I'm thinking I need to subtract the term on the RHS over to the LHS and do likewise with the y-term on the LHS over to the RHS so the powers of y are in an appropriate place.

OpenStudy (mendicant_bias):

\[\frac{dy}{dx}-\frac{y}{x}=-\frac{y^2}{x^2}\]

OpenStudy (mendicant_bias):

Where \[P(x)=-\frac{1}{x}, \ f(x) = -\frac{1}{x^2}, \text{and} \ n=2\]

OpenStudy (mendicant_bias):

The substitution made should be:\[u=y^{1-2}=y^{-1}=\frac{1}{y}\]

OpenStudy (anonymous):

Good so far.

OpenStudy (mendicant_bias):

Alright, so, taking care of the y'/adjusting for it being a function of different variables now:

OpenStudy (mendicant_bias):

\[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}= \ ? \]

OpenStudy (mendicant_bias):

(Having a brain fart, one moment)

OpenStudy (anonymous):

Moment granted :D

OpenStudy (mendicant_bias):

http://i.imgur.com/0IF6E2y.png

OpenStudy (anonymous):

Haha, nice.

OpenStudy (mendicant_bias):

Alright, actually legit forgetting how to think, sorry, could you....do the next step, lol?

OpenStudy (anonymous):

Okay, so \(u = y^{1-n}\), meaning we have \(u=y^{-1}\). Taking the derivative, we have: \(du = -y^{-2}y'\). Now we an multiply through by \(-y^{-2}\), giving us: \[-y^{-2}y' -y^{-2}(\frac{ - y}{ x }) = -y^{-2}(\frac{ -y^{2} }{ x^{2} })\] SInce \(-y^{-2}y' = du\) and \(y^{-1}\) = u, we can now make the appropriate substitutions: \[u' + \frac{ u }{ x } = \frac{ 1 }{ x^{2} }\]

OpenStudy (mendicant_bias):

Oh man, alright, reading.

OpenStudy (mendicant_bias):

Okay, so general procedure is once you've simplified and taken the derivative with respect to y of your substitution is to take everything in that derivative other than y prime itself and multiply through that value across the entire equation?

OpenStudy (anonymous):

Exactly. You wouldnt multiply through by the y', just the rest of it.

OpenStudy (mendicant_bias):

Okay, one sec.

OpenStudy (mendicant_bias):

Alright, I follow everything, how do we move forward from here? Is this essentially just....is using this technique on a Bernoulli DE just essentially transforming it into a form we can solve with already known, other techniques?

OpenStudy (anonymous):

Right, we transform it into a first order, linear ODE which is now solvable using an integrating factor.

OpenStudy (mendicant_bias):

:C I don't like integrating factors because of obvious reasons, well, alright, and from there, after we solve it with an integrating factor, do we just back-substitute in the values for our transformation in the first place?

OpenStudy (anonymous):

Yep. But yeah, as long as you know the process. If we skip the stupid steps, we have \[\mu = e^{\int\limits_{}^{}\frac{ 1 }{ x }dx} = x\] \[ux = \int\limits_{}^{}\frac{ 1 }{ x }dx\] \[ux = \ln|x| + C \implies y^{-1}x = \ln|x| + C\] And blah blah

OpenStudy (mendicant_bias):

(brb in about 5 minutes) Alright, cool

OpenStudy (mendicant_bias):

P.S. you still in that Starbucks?

OpenStudy (anonymous):

Nah, I went home early in a semi-emergency. Been home for a while now.

OpenStudy (mendicant_bias):

Ah. Back.

OpenStudy (mendicant_bias):

Alright, solved it as a Bernoulli to the point that we're at another technique, good enough for me. What I'm going to do now is try to solve it as a homogeneous DE but likewise not finish to solution, but just get it in the general form to be proven as a homogeneous DE, perform the necessary substitutions, and then leave it at that before moving on.

OpenStudy (anonymous):

sounds good.

OpenStudy (mendicant_bias):

Alright, taking the original setup of the problem: \[x^2 \frac{dy}{dx}+y^2=xy\]

OpenStudy (mendicant_bias):

\[x^2\frac{dy}{dx}=xy-y^2\]

OpenStudy (mendicant_bias):

\[\frac{dy}{dx}=\frac{xy-y^2}{x^2}\]

OpenStudy (mendicant_bias):

Nevermind, I went further than necessary, didn't I? Backtracking one second.

OpenStudy (anonymous):

I would do the substitution first as soon as I knew it was homogenous.

OpenStudy (mendicant_bias):

\[x^2 dy=(xy-y^2)dx\]

OpenStudy (mendicant_bias):

Well, putting it in the general form of a homogeneous DE sounds appropriate, yes? I just want to make sure I can get it in the right form first. And there is is, right?\[(xy-y^2)dx-x^2dy=0\]Homogeneous function of order II.

OpenStudy (mendicant_bias):

Alright, don't tell me how to do the next part, trying to remember how to proceed. Substitution like \[y = ux, \ \ \ x=vy\]

OpenStudy (anonymous):

Alright, not saying anything.

OpenStudy (mendicant_bias):

\[(xux-u^2x^2)dx-x^2(xdu+udx)=0\]

OpenStudy (mendicant_bias):

Is that general setup correct?

OpenStudy (anonymous):

Thats fine.

OpenStudy (mendicant_bias):

Alright, great, and then...from there, not going to do it yet, but conceptually, you do what again....never did it the first time, going to have to do it eventually.

OpenStudy (anonymous):

Now you do a normal separation of variables. Just the algebra to get there might suck a little.

OpenStudy (mendicant_bias):

"might suck a little" That sounds like this entire class, lol

OpenStudy (anonymous):

Aww, lol. I dont find too much of it annoying, only the series solutions stuff did I really not like.

OpenStudy (mendicant_bias):

Yeah, no, this stuff isn't all that bad, it's just that I didn't learn it right earlier and now hope I'm not going to send $1,000+ down the drain based on this final exam. It's high-stakes testing, and the immediate following morning I have another exam I am woefully unprepared for, and am having to invest all my time into this. Luckily, the other exam stuff is way easier to understand. Speaking of separation of variables, how the hell do I do that? I think we did a problem on it, give me a second to try to remember it myself.

OpenStudy (anonymous):

Let me know if you want me to mess with it :)

OpenStudy (mendicant_bias):

(The second exam, like this, is just an issue of time. ODE isn't really hard to understand....all of this stuff just takes time for consistency and performance.)

OpenStudy (mendicant_bias):

Alright, separation of variables is where you have a derivative term and two variables expressed in both present in the equation, you isolate all of them to one side and all of them to another and multiply through part of a differential, let's see.

OpenStudy (mendicant_bias):

One moment.

OpenStudy (mendicant_bias):

(Sorry lol)